Antuan Company set the following standard costs for one unit of its product. Direct materials (3.0 Ibs. @ $5.00 per Ib.) $ 15.00 Direct labor (1.8 hrs. @ $11.00 per hr.) 19.80 Overhead (1.8 hrs. @ $18.50 per hr.) 33.30 Total standard cost $ 68.10 The predetermined overhead rate ($18.50 per direct labor hour) is based on an expected volume of 75% of the factory’s capacity of 20,000 units per month. Following are the company’s budgeted overhead costs per month at the 75% capacity level. Overhead Budget (75% Capacity) Variable overhead costs Indirect materials $ 15,000 Indirect labor 75,000 Power 15,000 Repairs and maintenance 30,000 Total variable overhead costs $ 135,000 Fixed overhead costs Depreciation—Building 24,000 Depreciation—Machinery 71,000 Taxes and insurance 18,000 Supervision 251,500 Total fixed overhead costs 364,500 Total overhead costs $ 499,500 The company incurred the following actual costs when it operated at 75% of capacity in October. Direct materials (46,500 Ibs. @ $5.20 per lb.) $ 241,800 Direct labor (21,000 hrs. @ $11.20 per hr.)

235,200 Overhead costs Indirect materials $ 41,500 Indirect labor 176,900 Power 17,250 Repairs and maintenance 34,500 Depreciation—Building 24,000 Depreciation—Machinery 95,850 Taxes and insurance 16,200 Supervision 251,500 657,700 Total costs $ 1,134,700 3. Compute the direct materials cost variance, including its price and quantity variances. (Indicate the effect of each variance by selecting for favorable, unfavorable, and No variance.)
Answer #1

Direct material price variance = Actual quantity * (Standard price-Actual price)

Direct material price variance = 46,500 * ($5-5.20) = $9,300 Unfavorable

Direct material quantity variance = Standard price * (Standard quantity-Actual quantity)

Standard quantity = 15,000*3 = 45,000

Direct material quantity variance = $5 * (45,000-46,500) = $7,500 Unfavorable

Direct material cost variance = $7,500 U + 9,300 U = $16,800 Unfavorable